OpenStudy (anonymous):

If a baseball and bat cost $110, and the bat costs $100 more than the ball, how much does the ball cost?

OpenStudy (anonymous):

We can calculate this using a couple of equations. First, we need to define the two algebraically. We'll use A for the bat, and B for the ball. We know that both of these together cost $110. Therefore, A + B = 110. We also know that the bat costs $100 more than the ball. Therefore, A = B + 100. This is now our system of equations. A + B = 110 A = B + 100. Because we know the value of A as it relates to B, we can solve by substitution. We replace A with B + 100 to get our new equation. B + 100 + B = 110. Collect like terms: 2B + 100 = 110 Take away 100 from both sides: 2B = 10 Divide by two: B = 5. So now we know that the ball is worth 5 dollars. What about the bat? There are two ways we could do this. The first is that A + B = 110. A + 5 = 110. A = 105. The second way we can do this is that we know that the bat is worth 100 more than the ball. A = 100 + 5. A = 105. Both solutions come up the same: The bat is worth 105, and the ball is worth 5.